OpenStudy (anonymous):

Simplify, showing your steps. Tell the property you used in each step. 1/4 (1*4) + (-4+4)

OpenStudy (anonymous):

Is your question written like this one? \[\huge \frac{1}{4} (1 \times 4) + (-4+4)\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

If so then,

OpenStudy (anonymous):

\[\huge \frac{1}{4} (1 \times 4) + (-4+4)\] Do you know what is 1* 4

OpenStudy (anonymous):

and -4+ 4=?

OpenStudy (anonymous):

okay so 1/4(4)

OpenStudy (anonymous):

What properties would be used with each step?

OpenStudy (anonymous):

\[\huge \frac{1}{4} (1 \times 4) + (-4+4) \] \[\huge= \frac{1}{4} \times 4 + 0 = 1+ 0 = 1\]

OpenStudy (anonymous):

Okay so with each step which property did you use?

OpenStudy (anonymous):

1/4 * 4 = 1 and -4+4 = 0 so 1+0 = 1

OpenStudy (anonymous):

Alright then additive identity. Thank you